# Data Compression for Fast Online Stochastic Optimization

### v1

@@ Abstract
We propose an online data compression approach for efficiently solving distributionally robust optimization (DRO) problems with streaming data while maintaining out-of-sample performance guarantees. Our method dynamically constructs ambiguity sets using online clustering, allowing the clustered configuration to evolve over time for an accurate representation of the underlying distribution. We establish theoretical conditions for clustering algorithms to ensure robustness, and show that the performance gap between our online solution and the nominal DRO solution is controlled by the Wasserstein distance between the true and compressed distributions, which is approximated using empirical measures. We provide a regret analysis, proving that the upper bound on this performance gap converges sublinearly to a fixed clustering-dependent distance, even when nominal DRO has access, in hindsight, to the subsequent realization of the uncertainty. Numerical experiments in mixed-integer portfolio optimization demonstrate significant computational savings, with minimal loss in solution quality.
@@ 1 Introduction
Many decision-making problems in engineering, operations research, and computer science involve solving optimization problems affected by uncertain parameters. A common approach to address this class of problems is stochastic optimization (SO), which models uncertain parameters as random variables with known probability distributions. Typically, SO minimizes an expected cost based on this information [48]. While effective, this method relies on the assumption of complete knowledge of the underlying distributions—an assumption that is often unrealistic. If violated, SO produces suboptimal solutions that can lead to catastrophic consequences in the case of safety-critical systems. When faced with distributional ambiguity, i.e., uncertainty in the probability distribution, DRO provides a disciplined framework to robustify the decision-making process against distributional misspecification. Instead of assuming a known distribution, DRO considers an ambiguity set of possible distributions and optimizes for the worst-case expected cost under all distributions within this set. This approach is particularly appealing as it offers out-of-sample performance guarantees with a finite number of data samples. By accounting for ambiguity, DRO often outperforms non-robust methods such as sample average approximation (SAA) with respect to worst-case performance [22].
Since its appearance, significant effort has been devoted to constructing ambiguity sets primarily in static settings, where a single set is used throughout; see [16, 46, 40, 22], and the references therein. Even in multi-stage DRO, where uncertain data is revealed sequentially, a single ambiguity set is typically constructed, with time progression modeled through adjustable wait-and-see variables [18, 8, 6]. Fundamentally, this approach determines decisions based on a priori information. However, in many real-time applications—such as healthcare resource allocation, air traffic control, and energy trading—new data becomes available sequentially over time. As a result, the optimizer’s confidence in the nominal distribution can improve through the continuous assimilation of new data. Despite its potential, DRO with streaming data remains largely unexplored. In the few cases where it has been studied, updates to the DRO problem rely on gradient-based methods that either restrict the distribution to a finite set of scenarios [3], or require continuous decision variables and objective functions with local strong concavity [37]. In this work, we focus on the popular Wasserstein-based DRO [40, 36, 22], where the ambiguity set is defined as a Wasserstein ball around the empirical distribution of a dataset. As new data-points arrive, the problem dimension grows, increasing computational complexity [36, 56]. While a larger dataset improves confidence in the empirical measure as an approximation of the true distribution, the curse of dimensionality [21] ensures that the reduction in the ambiguity set’s radius is too slow to offset this complexity, even for moderately sized uncertain parameters. This leads to a computational bottleneck, and is especially limiting for real-time applications.
@@ 1.1 Our contributions
We propose an online data compression approach for efficiently solving Wasserstein DRO problems with streaming data, while maintaining out-of-sample performance guarantees. Our key contributions are:
  • Adaptive ambiguity sets via online clustering: We apply online clustering to construct adaptive ambiguity sets, formulated as Wasserstein balls of any order around a clustered empirical distribution. We establish theoretical conditions on clustering algorithms to ensure the robustness of our framework, which is compatible with any online clustering method that meets these conditions. We formalize the concept of optimal clustering, and provide a few fast and memory-efficient approximation algorithms suitable for our purposes.
  • Clustering effect analysis: We provide certificates, i.e., upper bounds on the out-of-sample performance of our algorithm, and prove that the performance gap between our compressed solution and the non-compressed DRO solution is controlled by the Wasserstein distance between the true distribution and its clustered approximation. This distance is approximated using empirical measures. We show that the objective converges to a value dependent on the number of clusters and the quality of the clustering algorithm. This quantifies the impact of data compression, and highlights the trade-off between computational effort and optimality.
  • Online performance analysis: We provide a regret analysis comparing the performance of our algorithm to the one of non-compressed DROin hindsight, i.e., with access to the subsequent realization of the uncertainty. We show that the regret converges sublinearly to a function of the Wasserstein distance between the true and compressed distributions, and with probability 1, the difference induced by the time-discrepancy converges to 0.
  • Computational gains in numerical experiments: We demonstrate the efficiency of our approach in sparse portfolio optimization. Our results show significant computational savings, even compared to SAA, with minimal loss in solution quality. We also demonstrate the possibility for the online framework to be memory-efficient, allowing the optimizer to discard data-points once seen, with minimal impact on solution times and quality.
@@ 1.2 Related work
Due to the ongoing data revolution, data stream clustering has recently attracted attention for emerging applications that involve large amounts of streaming data. They can be broadly classified into partition-based algorithms that partition data into clusters using distance-based similarity metrics, density-based algorithms that define clusters as dense partitions separated by sparse areas that dynamically change over time to adapt to data evolving distributions (such as DenStream [14]), hierarchical-based algorithms that maintain a tree-like structure by grouping similar clusters at different levels [42]. Among the first class, the most popular ones are undoubtedly incremental $k$-mean [1] and CluStream [2], which we adapt to fit our framework. We highlight that a major strength of our approach is its ability to incorporate state-of-the-art online clustering algorithms, with various memory and run-time complexities, with the choice left to the user’s discretion. In this work, we adapt two algorithms for our purposes.
@@ 1.3 Layout of the paper
In Section 2, we state the problem and introduce our online algorithm, and in Section 3, analyze the clustering effect by providing finite sample and asymptotic performance guarantees. In Section 4, we give dynamic regret bounds, and in Section 6, we describe various online clustering algorithms. In Section 7, we give guidelines for choosing the number of clusters and the ambiguity set radii experimentally, and in Section 8 we demonstrate our results on a portfolio optimization example.
@@ 2 The online stochastic problem
Consider a stochastic optimization problem of the form
@@
with each $f_{j}$ being proper, concave, and upper-semicontinuous in $u$ for all $x$. Additionally, we require the functions $f_{j}$ to be either Lipschitz or smooth for all $x\in\mathcal{X}$, according to the following definitions. Note that we assume the existence of the global constants by the compactness of $\mathcal{X}$.
Definition 2.1 (Lipschitzness).
A function $f(u,x)$ is $M$-Lipschitz on its domain with respect to the $\ell_{2}$-norm, if for all $x\in\mathcal{X}$,
$\left|f(v,x)-f(u,x)\right|\leq M\|u-v\|_{2},\quad\forall u,v\in\mathop{\bf dom}_{u}f.$
Definition 2.2 (Smoothness).
A differentiable function $f(u,x)$ is $L$-smooth on its domain, with respect to the $\ell_{2}$-norm, if for all $x\in\mathcal{X}$,
$\left\|\nabla f(v,x)-\nabla f(u,x)\right\|_{2}\leq L\|u-v\|_{2},\quad\forall u,v\in\mathop{\bf dom}_{u}f.$
Analogously to [56], we also assume $\mathop{\bf dom}_{u}{f}$ is independent of $x$, and satisfies the following assumption:
@@
The domain $\mathop{\bf dom}_{u}{f}$ is ${\mbox{\bf R}}^{d}$. Otherwise, $f$ is either element-wise monotonically increasing in $u$ and only has a (potentially) lower-bounded domain, or element-wise monotonically decreasing in $u$ and only has a (potentially) upper-bounded domain.
@@
We assume ${\mathbf{P}}$ to be unknown, and that the value $H_{\star}$ cannot be calculated as is. However, we have access to a streaming dataset of i.i.d. realization of $u$, which we use to construct an uncertainty framework. We begin with an initial dataset with $n_{0}\geq 1$ data-points, i.e., $\mathcal{D}_{0}=\{\hat{u}^{i}\}_{i=1}^{n_{0}}$. Then, at the end of each time step $t\geq 1$, the dataset is updated to be $\mathcal{D}_{t}=\{\hat{u}^{i}\}_{i=1}^{n_{0}+t}$, with $|\mathcal{D}_{t}|=n_{t}=n_{0}+t$. That is, we assume w.l.o.g. that one realization gets disclosed at the end of each round. Furthermore, we define the empirical distribution
$\hat{{\mathbf{P}}}_{t}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\delta_{\hat{u}^{i}},$
where $\delta_{u}$ is the Dirac distribution concentrating unit mass on $u$.
Distributionally robust optimization over time
[…]
$\mathcal{P}_{t}=\mathbf{B}_{\varepsilon_{t}}^{p}(\hat{{\mathbf{P}}}_{t})=\left\{\mathbf{Q}\in\mathcal{W}({S})\mid W_{p}(\hat{{\mathbf{P}}}_{t},\mathbf{Q})\leq\varepsilon_{t}\right\},$ (2)
@@
where the Wasserstein type-$p$ distance between two probability distributions $\mathbf{P},\mathbf{Q}\in\mathcal{W}(S)$ is defined by
$W_{p}(\mathbf{P},\mathbf{Q})=\inf_{\pi\in\Pi(\mathbf{P},\mathbf{Q})}\left(\int_{S\times S}\left\|u_{1}-u_{2}\right\|^{p}\mathrm{d}\pi\left(u_{1},u_{2}\right)\right)^{1/p},$ (3)
@@
and $\Pi(\mathbf{P},\mathbf{Q})$ is the set of all probability distributions over $S\times S$ with marginals $\mathbf{P}$ and $\mathbf{Q}$. Intuitively, the ambiguity set $\mathcal{P}_{t}$ contains all probability distribution in $\mathcal{W}(S)$ that can obtained by transporting probability mass from $\hat{{\mathbf{P}}}_{t}$ when the transportation budget is at most $\varepsilon_{t}$. With this uncertainty modeling framework in place, we compute the DRO solution $x_{t}\in\mathcal{X}$ at time step $t$ by solving the optimization problem
$H_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],$ (4)
where we seek to minimize the worst-case (w.r.t. any distribution in the ambiguity set) expected value of $f$. Note that we use the ambiguity set $\mathcal{P}_{t-1}$ constructed at time $t-1$: in the online framework, we assume to solve the problem before the new data-point is revealed.
As $t$ increases, our confidence in $\hat{{\mathbf{P}}}_{t}$ grows, allowing us to safely reduce the radius $\varepsilon_{t}$. This radius can be interpreted as a measure of trust in the data, and the cost of robustness. This intuition is supported by the results in [20], which show that under mild assumptions, choosing $\varepsilon_{t}\sim\text{o}(t^{-\frac{1}{d}})$ ensures that the true distribution $\mathbf{P}\in\mathcal{P}_{t}$ with high probability. From [40], this choice of radius leads to the following finite sample performance guarantee
@@
${\mathbf{P}}^{t-1}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x_{t})]\leq H_{t}\right)\geq 1-\beta_{t-1},$ (5)
where for all $t\geq 1$, $\beta_{t-1}>0$ is a specified time-varying probability of constraint violation, and ${\mathbf{P}}^{t-1}$ is the product distribution of the dataset $\mathcal{D}_{t-1}$. The in-sample objective value $H_{t}$ is then a certificate for the out-of-sample performance ${\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x_{t})]$. Furthermore, [40] prove that an appropriately decreasing choice of $\varepsilon_{t}$ and $\beta_{t}$ as $t\to\infty$ leads to asymptotic guarantees on the convergence of $H_{t}$ to the true value $H_{\star}$. These results imply, by incorporating streaming data into Problem (4), we can adapt the ambiguity set and gradually reduce the cost of robustness.
Our online algorithm

### v2

# Fast Online Distributionally Robust Optimization via Data Compression

@@ Abstract
We propose an online data compression approach for efficiently solving distributionally robust optimization (DRO) problems with streaming data while maintaining out-of-sample performance guarantees. Our method dynamically constructs ambiguity sets using online clustering, allowing the clustered configuration to evolve over time for an accurate representation of the underlying distribution. We establish theoretical conditions for clustering algorithms to ensure robustness, and show that the performance gap between our online solution and the nominal DRO solution can be written in terms of the distance between the true and compressed distributions. Therefore, by varying the number of clusters, our method effectively balances robustness and online computational efficiency. We show that our analysis is compatible with well-established finite-sample and asymptotic guarantees for Wasserstein DRO. Numerical experiments in mixed-integer portfolio optimization demonstrate significant computational savings, with minimal loss in solution quality.
@@ 1 Introduction
Many decision-making problems in engineering, operations research, and computer science involve solving optimization problems affected by uncertain parameters. Distributionally robust optimization (DRO) has emerged as a disciplined framework to robustify the decision-making process against distributional misspecification [43, 34]. Instead of assuming a known distribution, DRO considers an ambiguity set of possible distributions and optimizes for the worst-case expected cost under all distributions within this set. This approach is particularly appealing as it offers out-of-sample performance guarantees with a finite number of data samples. By accounting for ambiguity, DRO often outperforms non-robust methods such as sample average approximation (SAA), especially with respect to worst-case performance [18].
Since its appearance, significant effort has been devoted to constructing ambiguity sets primarily in static settings, using either structural assumptions on the distribution (e.g., moments [44, 21]) or training data directly [33, 18]. However, many real-time settings, such as healthcare resource allocation, air traffic control, and energy trading, are dynamic and new data becomes available sequentially over time. Despite its potential, DRO has been largely underdeveloped in such scenarios. Most DRO approaches construct a static ambiguity set, and model time progression through adjustable wait-and-see variables [15, 8, 6]. There are limited works focusing on how to adapt the ambiguity sets dynamically over time, where they pose strict assumptions on the problem: either by restricting the distribution to a finite set of scenarios [3], or by requiring continuous variables and objective functions with local strong concavity [35]. In particular, streaming data settings can be challenging for the popular Wasserstein-based DRO [38, 33, 18], where the ambiguity set is defined as a Wasserstein ball around the empirical distribution of a dataset. As new datapoints arrive, the problem dimension grows, increasing computational complexity [33, 54]. While a larger dataset improves confidence in the empirical measure as an approximation of the true distribution, the reduction in the ambiguity set’s radius is not enough to offset this complexity; for mixed-integer problems, even SAA is difficult to solve with large amounts of data, without any added complexity from ambiguity. This leads to computational bottlenecks that are especially limiting for real-time applications.
@@ 1.1 Our contributions
We propose an online data compression approach for efficiently solving Wasserstein DRO problems with streaming data, while also maintaining out-of-sample performance guarantees. We allow for mixed-integer formulations, and a broad class of underlying distributions. Our key contributions are:
  • Adaptive ambiguity sets via online clustering: We apply online clustering to construct adaptive ambiguity sets, formulated as Wasserstein balls of any order around a clustered empirical distribution. We establish theoretical conditions on clustering algorithms to ensure the robustness of our framework, which is compatible with any online clustering method that meets these conditions. We formalize the concept of optimal clustering, and provide three fast and memory-efficient approximation algorithms suitable for our purposes.
  • Clustering effect analysis: We prove that the performance gap between our compressed solution and the non-compressed (nominal) DRO solution can be written in terms of the distance between the true distribution and its clustered approximation, and the ambiguity set radius. Under mild assumptions, this radius can decay at a rate independent of the uncertainty dimension. We also show that our objective converges to a value dependent on the number of clusters and the quality of the clustering algorithm. This quantifies the impact of data compression, and highlights the trade-off between computational effort and optimality. We provide both finite-sample and asymptotic performance guarantees for our solution.
  • Computational gains in numerical experiments: We demonstrate the efficiency of our approach in sparse portfolio optimization. Our results show significant computational savings, even compared to SAA, with minimal loss in solution quality. We also demonstrate the possibility for the online framework to be memory-efficient, allowing the optimizer to discard datapoints once seen, with minimal impact on solution times and quality.
@@ 1.2 Related work
Due to the ongoing data revolution, data stream clustering has recently attracted attention for emerging applications that involve large amounts of streaming data. They can be broadly classified into partition-based algorithms that partition data into clusters using distance-based similarity metrics, density-based algorithms that define clusters as dense partitions separated by sparse areas that dynamically change over time to adapt to data evolving distributions (such as DenStream [12]), hierarchical-based algorithms that maintain a tree-like structure by grouping similar clusters at different levels [40]. Among the first class, the most popular ones are undoubtedly incremental $k$-mean [1] and CluStream [2], which we adapt to fit our framework. We point to the survey [13] for further details. We highlight that a major strength of our approach is its ability to incorporate state-of-the-art online clustering algorithms, with various memory and run-time complexities, with the choice left to the user’s discretion.
@@ 1.3 Layout of the paper
In Section 2, we state the problem and introduce our online algorithm, and in Section 3, analyze the clustering effect by providing finite sample and asymptotic performance guarantees. In Section 4, we give theoretical results for the specific case of bounded support. In Section 5, we describe various online clustering algorithms, and in Section 6, give guidelines for choosing the number of clusters and the ambiguity set radii experimentally. In Section 7 we demonstrate our results on a portfolio optimization example.
@@ 2 The online stochastic problem
We consider a stochastic optimization (SO) problem of the form
@@
with each $f_{j}$ being proper, concave, and upper-semicontinuous in $u$ for all $x$. Additionally, we require the functions $f_{j}$ to be either Lipschitz or smooth for all $x\in\mathcal{X}$, according to the following definitions. We assume the existence of the global constants independent of $x$ by the compactness of $\mathcal{X}$.
Assumption 2.1 (Lipschitzness and smoothness).
Let at least one hold.
  1. 1.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ are Lipschitz continuous with constants $M_{j}$, i.e., $\left|f_{j}(v,x)-f_{j}(u,x)\right|\leq M_{j}\|u-v\|_{2},\;\forall u,v\in\mathop{\bf dom}_{u}f$.
  2. 2.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ are smooth with constants $L_{j}$, i.e., $\left\|\nabla f_{j}(v,x)-\nabla f_{j}(u,x)\right\|_{2}\leq L_{j}\|u-v\|_{2},\;\forall u,v\in\mathop{\bf dom}_{u}f$.
  3. 3.
    For all $x\in\mathcal{X}$, $f$ is smooth with constant  $L$, i.e., $\left\|\nabla f(v,x)-\nabla f(u,x)\right\|_{2}\leq L\|u-v\|_{2},\;\forall u,v\in\mathop{\bf dom}_{u}f$.
Analogously to [54], we assume the following on the domain of $f$.
@@
The domain $\mathop{\bf dom}_{u}{f}$ is ${\bf R}^{d}$. Otherwise, $f$ is either element-wise monotonically increasing in $u$ and only has a (potentially) lower-bounded domain, or element-wise monotonically decreasing in $u$ and only has a (potentially) upper-bounded domain. In all cases, the domain $\mathop{\bf dom}_{u}{f}$ is independent of $x$.
@@
We assume the distribution $\mathbf{P}$} to be unknown, and that the value $H_{\star}$ cannot be calculated as is. However, we have access to a streaming dataset of i.i.d. realization of $u$, which we use to construct an uncertainty framework. We begin with an initial dataset with $n_{0}\geq 1$ datapoints, i.e., $\mathcal{D}_{0}=\{\hat{u}^{i}\}_{i=1}^{n_{0}}$. Then, at the end of each time step $t\geq 1$, the dataset is updated to be $\mathcal{D}_{t}=\{\hat{u}^{i}\}_{i=1}^{n_{0}+t}$, with $|\mathcal{D}_{t}|=n_{t}=n_{0}+t$. That is, we assume w.l.o.g. that one realization gets disclosed at the end of each time step. Furthermore, we define the empirical distribution $\hat{{\mathbf{P}}}_{t}=(1/{n_{t}})\sum_{i=1}^{n_{t}}\delta_{\hat{u}^{i}},$ where $\delta_{u}$ is the Dirac distribution concentrating unit mass on $u$.
2.1 DRO over time
Let $\Xi_{p}(S)$ be the space of all probability distributions $\mathbf{Q}$ defined on the support $S$, with bounded $p$-th moments, i.e., ${\mbox{\bf E}}_{\mathbf{Q}}[\|u\|^{p}]=\int_{S}\|u\|^{p}\mathbf{Q}(du)<\infty$, where $\|\cdot\|$ is an arbitrary norm on ${\bf R}^{d}$, and either $p\geq 1$ or $p=\infty$.
The DRO approach under the Wasserstein metric constructs an ambiguity set of order $r$ around $\hat{{\mathbf{P}}}_{t}$ of radius $\varepsilon_{t}\geq 0$, i.e.,
$\mathcal{P}_{t}=\mathbf{B}_{\varepsilon_{t}}^{r}(\hat{{\mathbf{P}}}_{t})=\left\{\mathbf{Q}\in\Xi_{r}({S})\mid W_{r}(\hat{{\mathbf{P}}}_{t},\mathbf{Q})\leq\varepsilon_{t}\right\},$ (2)
@@ 2.1 DRO over time
where the Wasserstein type-$p$ distance between two probability distributions $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\Xi_{p}(S)$ is defined by
$W_{p}(\mathbf{Q}_{1},\mathbf{Q}_{2})=\inf_{\pi\in\Pi(\mathbf{Q}_{1},\mathbf{Q}_{2})}\left(\int_{S\times S}\left\|u_{1}-u_{2}\right\|^{p}\mathrm{d}\pi\left(u_{1},u_{2}\right)\right)^{1/{p}},$ (3)
and $\Pi(\mathbf{Q}_{1},\mathbf{Q}_{2})$ is the set of all probability distributions over $S\times S$ with marginals $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$. In this work, we always refer to $r$ as the order of the Wasserstein distance used within an ambiguity set. For other Wasserstein distances used in the analysis, we refer to order $p$. In addition, whenever we use a distance metric of order $p$ on a distribution $\mathbf{Q}$, we assume that the distribution itself has finite $p$-th moments. We suppress the dependency of $\mathcal{P}_{t}$ on $r$ for conciseness, and note that unless otherwise specified, $r$ is of an arbitrary order and $\|\cdot\|$ is an arbitrary norm. Intuitively, the ambiguity set $\mathcal{P}_{t}$ contains all probability distribution in $\Xi(S)$ that can obtained by transporting probability mass from $\hat{{\mathbf{P}}}_{t}$ when the transportation budget is at most $\varepsilon_{t}$.
@@
With this uncertainty modeling framework in place, we compute the DRO solution $x_{t}\in\mathcal{X}$ at time step $t$ by solving the optimization problem
$H_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],$ (4)
where we seek to minimize the worst-case (w.r.t. any distribution in the ambiguity set) expected value of $f$.
2.2 Performance guarantees
As $t$ increases, our confidence in $\hat{{\mathbf{P}}}_{t}$ increases, allowing us to safely reduce the radius $\varepsilon_{t}$, which can be interpreted as a measure of trust in the data. With suitable choices of $\varepsilon_{t}$ [17, 38, 18], which we will refer to as the nominal radius, we obtain finite sample performance guarantees of the form
@@ 2.2 Performance guarantees
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{P}}^{t}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x_{t})]\leq H_{t}+\rho_{t}\right)\geq 1-\beta_{t},$} (5)
where for all $t\geq 1$, $\beta_{t}>0$ is a specified time-varying probability of constraint violation, $\rho_{t}$ is a residual term, and ${\mathbf{P}}^{t}$ is the product distribution of the dataset $\mathcal{D}_{t}$. In this work, we make use of two such choices of $\varepsilon_{t}$, which require the distribution $\mathbf{P}$} to satisfy distinct assumptions. We summarize them here.
Dimension-dependent
One choice is to impose a light-tailed assumption, which is satisfied by sub-Gaussian distributions and by bounded distributions.
Assumption 2.3 (Light-tailed distribution).
There exists an exponent $\alpha>1$, such that $R={\mbox{\bf E}}_{\mathbf{P}}[\exp(\|u\|^{\alpha})]=\int_{S}\exp(\|u\|^{\alpha}){\mathbf{P}}(du)<\infty.$
Then, measure concentration results for Wasserstein-$p$ distances imply the following.
Theorem 2.1 (Guarantees using Wasserstein measure concentration [17, 33]).
Let $p$ be an arbitrary order. Under Assumption 2.3 for the case $p\geq 1$, and under analogous assumptions for $p=\infty$ [47], for $n_{t}\rightarrow\infty$ we can choose confidence levels $\beta_{t}\in(0,1)$ and corresponding radii $\varepsilon_{t}=O((\log(\beta^{-1}_{t})/{n_{t}})^{\min(p/d,1/2)})$ such that $\sum_{t=1}^{\infty}\beta_{t}<\infty$, ${\mathbf{P}}^{t}(W_{p}({\mathbf{P}},\hat{{\mathbf{P}}}_{t})\geq\varepsilon_{t})\leq\beta_{t},$ and $\lim_{t\rightarrow\infty}\varepsilon_{t}=0.$ The DRO problem (4) solved with $r=p$ and these radii gives the finite-sample guarantee (5) with $\rho_{t}=0$.
We point to the cited works [17, 47, 33] for the constants involved. In Section 4, we further calculate explicit radii for the specific case of bounded support $S$. From the above characterization, it is evident that larger datasets are beneficial for reducing the distributional ambiguity around the nominal distribution. Indeed, we can use Theorem 2.1 to conclude that $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W_{r}(\mathbf{P},\hat{\mathbf{P}}_{t}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=0}\}=1$, which implies $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}H_{t}\downarrow H_{\star}\}=1$ under the Lipschitz condition of Assumption 2.1 [38, Theorem 3.7]. However, as the radius scales with $O(n_{t}^{-r/d})$, even for moderately sized $d$ this bound suffers from the curse of dimensionality. The slow decrease in the required radius while the dataset grows makes the resolution of the DRO problem in (4) computationally challenging, as a constraint needs to be added for each datapoint constituting the empirical distribution $\hat{\mathbf{P}}_{t}$ (see formulation (9)). Clearly, for $t\rightarrow\infty$, we would end up with an infinite-dimensional problem that cannot be solved.
Dimension-free
Recent work [18] decouples the radii $\varepsilon_{t}$ from the dimension $d$ by providing finite sample guarantees for type-$r$ Wasserstein DRO, $r\in[1,2]$, with radii scaling as $\tilde{O}(n_{t}^{-1/2})$ ($\tilde{O}$ suppresses logarithmic terms). These result require $\mathbf{P}$} to satisfy a transportation-information inequality, and $f$ to satisfy certain curvature conditions, stated below for completeness.
Assumption 2.4.
If we consider $r=1$ for (2), then $\mathbf{P}$} satisfies the transportation-information inequality $T_{1}(\alpha)$ for some positive constant $\alpha$, i.e., for all $\mathbf{Q}\in\Xi_{1}(S)$, $W_{1}(\mathbf{Q},{\mathbf{P}})\leq\sqrt{\alpha H(\mathbf{Q}\|{\mathbf{P}})}$, where $H(\mathbf{Q}\|{\mathbf{P}})$ denotes the relative entropy. Furthermore, $f$ satisfies clause 1 of Assumption 2.1. If $r=2$, then $\mathbf{P}$} satisfies $T_{2}(\alpha)$ for some positive constant $\alpha$, and $f$ satisfies clause 3 of Assumption 2.1.
If the objective function class $\mathcal{F}=\{f(\cdot,x)\mid x\in\mathcal{X}\}$ additionally satisfies certain assumptions on its Rademacher complexity [18, Assumption 5], the following holds.
Theorem 2.2 (Dimension-free performance guarantees [18]).
For orders $r=1$ and $r=2$, under Assumption 2.4 and a Rademacher complexity condition [18, Assumption 5], we can choose confidence levels $\beta_{t}=O(\log(n_{t}))\in(0,1)$, radii $\varepsilon_{t}=\tilde{O}(n_{t}^{-1/2})$, and residual terms $\rho_{t}=\tilde{O}(1/n_{t})$ such that for all $x\in\mathcal{X}$,
${\mathbf{P}}^{t}\left({\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x)]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x)]+\rho_{t}\right)\geq 1-\beta_{t}.$ (6)
This implies the finite-sample guarantee (5) with the above parameter values.
In exchange for dimension-free radii $\varepsilon_{t}$, this result introduces a nonzero residual $\rho_{t}$. However, even with smaller radii, the difficulty of solving the DRO problem with a large number of datapoints can be high. We are thus motivated to introduce an additional tradeoff, where a separate residual term is established in exchange for a more compact DRO formulation. This leads to our framework below.
2.3 Our online algorithm